RC Meeting: September 4th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings


2020

March
19, 20, 21, 22, 23, 24
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

RC:

Summary of Sep. 4th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3840690
  2. Issues: High rates on upstream halo counters; only remedy proposed by MCC is raising the FSD threshold to 2600.
  3. Recorded data: 32.2 M events

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3840902
  2. Issues: No beam - accelerator trying to tune orbits, then 1L19 drops out again. Attempt to tune beam on tagger failed (high upstream counter rates even with beam to Hall off).
  3. Recorded data: 0

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3841144
  2. Issues: Still no beam. MCAT04 magnet may have been energized to low (would explain scraping): https://logbooks.jlab.org/entry/3841133. Finally beam at the very end, but DAQ troubles and then cold box crash.
  3. Recorded data: 3.2 M events

Whole day summary

  1. Total recorded data today: 35.4 M events - 1 good shift's worth in 3 shifts
  2. Part-II total accumulated data on D2: 1.54 B events
  3. Concerns: Still unsolved: Occasionally high upstream counters, vacuum 2H01, lots of beam crashes and instability


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded